# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ Appendix C Extension to U-statistics
  • Extending the qualitative (non-)robustness results. The main ingredients for proving Theorem 1 are (i) a deviation bound for the V-statistic $D^{2}(\mathbb{X}_{n})$ as shown in Lemma 7, and (ii) a concentration bound for the bootstrapped quantile $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})$ as shown in Lemma 8. A U-statistic counterpart of both results can be shown by following the same proof technique and replacing $D^{2}(\mathbb{X}_{n})$ with the U-statistic. For example, with the U-statistic, a decomposition similar to (10) can be derived by summing over only the non-diagonal terms $i\neq j$ and replacing normalizing factors of the form $1/n^{2}$ to $1/(n(n-1))$. Similarly, the proof of Theorem 3 can be extended to U-statistics by following the same steps in Section A.3.
@@ Appendix D A Non-Asymptotically Valid Robust KSD Test
The robust-KSD test introduced in Section 4 is only well-calibrated when $n\rightarrow\infty$. In this appendix, we now derive a robust test that is well-calibrated with finite samples. The test rejects the null $H^{\mathrm{C}}_{0}$ in (8) if
@@
where $\gamma_{n}=\sqrt{\tau_{\infty}/n}+\sqrt{-2\tau_{\infty}(\log\alpha)/n}$, the constant $\tau_{\infty}$ is defined in Lemma 2, and $\Delta_{\theta}(\mathbb{X}_{n})$ is defined in (9). We call this test robust-KSD-Dev (R-KSD-Dev).
The decision threshold $\gamma_{n}$ is based on the concentration bound Lemma 18, which is a deviation bound of the McDiarmid’s type (McDiarmid et al.,, 1989). This test can be viewed as a counterpart of the robust MMD test of Sun and Zou, (2023), which is also constructed using McDiarmid’s inequality. The next result, proved at the end of this section, shows its finite-sample validity under $H^{\mathrm{C}}_{0}$ as well as its consistency under $H^{\mathrm{C}}_{1}$.
@@
Let $\mathbb{X}_{\infty}=\{\mathbf{X}_{i}\}_{i=1}^{\infty}$ be a sequence of independent random variables following $Q$. Suppose $k$ is a tilted kernel satisfying the conditions in Lemma 2, and let $\theta\geq 0$. Then
@@
Alternative deviation bounds to Lemma 18 can also be used to construct similar tests. More precisely, Theorem 24 holds for any threshold $\gamma_{n}$ that satisfies
@@
where $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\big{\|}\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{Y})]\big{\|}_{\mathcal{H}_{u}}$ is a Hilbert-space norm, as shown in Lemma 16. Thus, any deviation bound for Hilbert-space norms may be applied. Examples include another McDiarmid’s bound of Gretton et al., (2012, Theorem 7), the empirical Bernstein bounds of Wolfer and Alquier, (2022, Theorem A.1) and Martinez-Taboada and Ramdas, (2024, Corollary 1), as well as the Hilbert-space valued Hoeffding bound of Pinelis, (1994, Theorem 3.5), particularly its i.i.d. variant (Chatalic et al.,, 2022, Lemma E.1).
@@
We opted for the McDiarmid-type bound in Lemma 18 because, in our setting, it is tighter than the bounds from Gretton et al., (2012), Wolfer and Alquier, (2022) and Chatalic et al., (2022). In particular, while Wolfer and Alquier, (2022) claim their empirical Bernstein bound outperforms McDiarmid’s inequality, we find that this only holds for their bound tailored to translation-invariant kernels (Wolfer and Alquier,, 2022, Theorem 3.1), but not for the non-translation-invariant version (Wolfer and Alquier,, 2022, Theorem A.1), which our setting requires since we have assumed Lemma 2. Moreover, although the bound in Lemma 18 is worse than the empirical Bernstein bound of Martinez-Taboada and Ramdas, (2024), we find that the difference is only marginal, and the former is considerably easier to implement.
In Figure 15, we compare R-KSD-Dev with the bootstrap-based robust-KSD test, the robust MMD test and the standard tests under the outlier-contaminated Gaussian model in Section 5.1. We set the uncertainty radius to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ for various values of $\epsilon_{0}$, where $\tau_{\infty}$ is estimated following Section 4.2. As expected, R-KSD-Dev controls Type-I error against Huber’s contamination with maximal proportion $\epsilon_{0}$. However, when $\epsilon>\epsilon_{0}$, it has lower power compared with robust-KSD and the MMD-based tests. This is because R-KSD-Dev uses a deviation bound to construct the decision threshold, which is a uniform bound over all data-generating distributions and is therefore conservative. For the same reason, we expect tests constructed using the alternative deviation bounds discussed in Remark 24.1 to suffer from similar conservativeness.
@@
Figure 15: Rejection probability under an outlier-contaminated Gaussian model. Black dash-dot vertical line. Uncertainty radius $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, with $\tau_{\infty}$ estimated following Section 4.2.
Proof of Theorem 24  Proceeding as in the proof of Theorem 4, for any $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, we have the inequality
@@
By the deviation bound in Lemma 18, the RHS is bounded by $\alpha$, thus proving the first claim.
@@
The same argument in Section A.5.2 shows that $\sqrt{n}\big{(}D(\mathbb{X}_{n})-D(Q,P)\big{)}$ converges weakly to a Gaussian distribution. On the other hand, since $D(Q,P)>\theta$ under $H^{\mathrm{C}}_{1}$ and $\gamma_{n}\to 0$ as $n\to\infty$, the term $\sqrt{n}\big{(}\gamma_{n}+\theta-D(Q,P)\big{)}\to-\infty$. Therefore, the probability in the last line converges to one, thus proving the second claim.

### v4

@@ Appendix C Extension to U-statistics
  • Extending the qualitative (non-)robustness results. The main ingredients for proving Theorem˜1 are (i) a deviation bound for the V-statistic $D^{2}(\mathbb{X}_{n})$ as shown in Lemma˜7, and (ii) a concentration bound for the bootstrapped quantile $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})$ as shown in Lemma˜8. A U-statistic counterpart of both results can be shown by following the same proof technique and replacing $D^{2}(\mathbb{X}_{n})$ with the U-statistic. For example, with the U-statistic, a decomposition similar to (10) can be derived by summing over only the non-diagonal terms $i\neq j$ and replacing normalizing factors of the form $1/n^{2}$ to 1/(n(n1))1/(n(n-1)). Similarly, the proof of Theorem˜3 can be extended to U-statistics by following the same steps in Section˜A.3.
@@ Appendix D A Non-Asymptotically Valid Robust KSD Test
The robust-KSD test introduced in Section˜4 is only well-calibrated when $n\rightarrow\infty$. In this appendix, we now derive a robust test that is well-calibrated with finite samples. An immediate consequence is that a stronger, uniform Type-I error control can be achieved.
The test rejects the null $H^{\mathrm{C}}_{0}$ in (8) if
@@
where $\gamma_{n}=\sqrt{\tau_{\infty}/n}+\sqrt{-2\tau_{\infty}(\log\alpha)/n}$, the constant $\tau_{\infty}$ is defined in Lemma˜2, and $\Delta_{\theta}(\mathbb{X}_{n})$ is defined in (9). We call this test robust-KSD-Dev (R-KSD-Dev).
The decision threshold $\gamma_{n}$ is based on the concentration bound Lemma˜18, which is a deviation bound of the McDiarmid’s type (McDiarmid et al.,, 1989). This test can be viewed as a counterpart of the robust MMD test of Sun and Zou, (2023), which is also constructed using McDiarmid’s inequality. The next result, proved at the end of this section, shows its finite-sample validity under $H^{\mathrm{C}}_{0}$ as well as its consistency under $H^{\mathrm{C}}_{1}$.
@@
Let $\mathbb{X}_{\infty}=\{\mathbf{X}_{i}\}_{i=1}^{\infty}$ be a sequence of independent random variables following $Q$. Suppose $k$ is a tilted kernel satisfying the conditions in Lemma˜2, and let $\theta\geq 0$. Then
@@
Alternative deviation bounds to Lemma˜18 can also be used to construct similar tests. More precisely, Theorem˜24 holds for any threshold $\gamma_{n}$ that satisfies
@@
where $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\big{\|}\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{Y})]\big{\|}_{\mathcal{H}_{u}}$ is a Hilbert-space norm, as shown in Lemma˜16. Thus, any deviation bound for Hilbert-space norms may be applied. Examples include another McDiarmid’s bound of Gretton et al., (2012, Theorem 7), the empirical Bernstein bounds of Wolfer and Alquier, (2022, Theorem A.1) and Martinez-Taboada and Ramdas, (2024, Corollary 1), as well as the Hilbert-space valued Hoeffding bound of Pinelis, (1994, Theorem 3.5), particularly its i.i.d. variant (Chatalic et al.,, 2022, Lemma E.1).
@@
We opted for the McDiarmid-type bound in Lemma˜18 because, in our setting, it is tighter than the bounds from Gretton et al., (2012), Wolfer and Alquier, (2022) and Chatalic et al., (2022). In particular, while Wolfer and Alquier, (2022) claim their empirical Bernstein bound outperforms McDiarmid’s inequality, we find that this only holds for their bound tailored to translation-invariant kernels (Wolfer and Alquier,, 2022, Theorem 3.1), but not for the non-translation-invariant version (Wolfer and Alquier,, 2022, Theorem A.1), which our setting requires since we have assumed Lemma˜2. Moreover, although the bound in Lemma˜18 is worse than the empirical Bernstein bound of Martinez-Taboada and Ramdas, (2024), we find that the difference is only marginal, and the former is considerably easier to implement.
In Figure˜15, we compare R-KSD-Dev with the bootstrap-based robust-KSD test, the robust MMD test and the standard tests under the outlier-contaminated Gaussian model in Section˜5.1. We set the uncertainty radius to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ for various values of $\epsilon_{0}$, where $\tau_{\infty}$ is estimated following Section˜4.2. As expected, R-KSD-Dev controls Type-I error against Huber’s contamination with maximal proportion $\epsilon_{0}$. However, when $\epsilon>\epsilon_{0}$, it has lower power compared with robust-KSD and the MMD-based tests. This is because R-KSD-Dev uses a deviation bound to construct the decision threshold, which is a uniform bound over all data-generating distributions and is therefore conservative. For the same reason, we expect tests constructed using the alternative deviation bounds discussed in Remark˜24.1 to suffer from similar conservativeness.
@@
Figure 15: Rejection probability under an outlier-contaminated Gaussian model. Black dash-dot vertical line. Uncertainty radius $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, with $\tau_{\infty}$ estimated following Section˜4.2.
Proof of Theorem˜24  Proceeding as in the proof of Theorem˜4, for any $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, we have the inequality
@@
By the deviation bound in Lemma˜18, the RHS is bounded by $\alpha$, thus proving the first claim.
@@
The same argument in Section˜A.5.2 shows that $\sqrt{n}\big{(}D(\mathbb{X}_{n})-D(Q,P)\big{)}$ converges weakly to a Gaussian distribution. On the other hand, since $D(Q,P)>\theta$ under $H^{\mathrm{C}}_{1}$ and $\gamma_{n}\to 0$ as $n\to\infty$, the term $\sqrt{n}\big{(}\gamma_{n}+\theta-D(Q,P)\big{)}\to$-$\infty$. Therefore, the probability in the last line converges to one, thus proving the second claim.